OpenStudy (anonymous):

What is the area of a trapezoid that has bases of 1feet and 14 inches and a height of 3 inches?

OpenStudy (anonymous):

area of a trapezoid is.. A=[(b1+b2)*h]/2 1 foot into inches is 12 making the equation [(12+14)*3]/2. A=[(26)*3]/2 A=78/2 A=39in^3. Hope this helps :)

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

Your welcome!

OpenStudy (anonymous):

SO WHERE YOUR FROM